# skipabove=1em,skipbelow=1em

### v3

@@ Abstract.
Figure 1. Overall framework of ingestion health sensing and intervention
Ingestive behavior plays a critical role in health, yet many existing interventions remain limited to static guidance or manual self-tracking. With the increasing integration of sensors and perceptual computing, recent systems have begun to support closed-loop interventions that dynamically sense user behavior and provide feedback during or around ingestion episodes. In this survey, we review 136 studies that leverage sensor-enabled or interaction-mediated approaches to influence eating behavior. We propose a behavioral closed-loop paradigm comprising three core components: target behaviors, sensing modalities, and feedback strategies. A taxonomy of sensing and intervention modalities is presented, organized along human- and environment-based dimensions. Our analysis also examines evaluation methods and design trends across different modality-behavior pairings. This review reveals prevailing patterns and critical gaps, offering design insights for future adaptive and context-aware ingestion health interventions.
@@ 1. Introduction
Ingestion health has been shown to significantly influence quality of life, longevity, and overall well-being (Jovanov et al., 2014; Mesas et al., 2012). Given its critical importance, it is essential to develop effective strategies that integrate both sensing and intervention for ingestive behaviors. Traditional approaches to ingestion monitoring and intervention, such as food diaries (Thompson and Subar, 2017), self-report questionnaires (Shim et al., 2014), and periodic counseling sessions (Greaves et al., 2011), are heavily dependent on user self-discipline and engagement. Moreover, these methods are typically static, offering one-time or episodic interventions without continuous feedback mechanisms (Raber et al., 2021; Takemoto et al., 2018). As a result, they often fail to dynamically support individuals in developing sustainable and healthy ingestive behaviors over time.
With the advancement of sensing technologies, new opportunities have emerged for low-burden and continuous sensing of ingestive behaviors. These technologies not only support the development of systems capable of capturing rich behavioral data with minimal user effort, but also enable more personalized and adaptive feedback and intervention strategies (Bell et al., 2020). Consequently, there is a growing body of research that leverages sensing and interactive technologies to bridge the gap between data collection and real-time intervention in ingestion health. This shift represents a critical evolution from static, user-driven methods toward context-aware, behavior-informed support systems.
Building on these advances, research in behavioral sensing and intervention has expanded considerably in recent years, alongside growing interest in ingestion health. Several recent reviews have addressed related domains, which can be broadly categorized into three research streams: (1) dietary intake and nutrition assessment, (2) wearable ingestion detection technologies, and (3) behavior change interventions.
Dietary Intake Nutrition Assessment
Traditional dietary assessment methods face challenges in both accuracy and user burden, prompting researchers to explore sensor-based technologies as a more objective alternative. Skinner et al. (2020) proposed a framework that integrates physical, chemical, and biosensors to objectively capture what, how, and when individuals eat. Wang et al. (2022) reviewed real-time sensor systems for nutritional support, while emphasizing deployment challenges. While these advances improve the reliability of intake measurement objectivity, they remain largely confined to assessment purposes and lack extensions toward adaptive behavioral interventions in ingestion health.
Wearable Ingestion Detection Technologies
Wearable-based ingestion detection has been extensively reviewed from different perspectives. Bell et al. (2020) surveyed automatic intake detection studies in free-living settings, highlighting sensor modalities and validation metrics. He et al. (2020) categorized detection systems by sensor type, while Hiraguchi et al. (2023) provided a comprehensive survey covering both eating and drinking behaviors. Beyond general ingestion detection, specialized methods for chewing detection have been explored. Wei et al. (2022) and Mortazavi and Gutierrez-Osuna (2023) reviewed wearable-based chewing sensing techniques and signal processing approaches. Despite these efforts, existing studies and reviews consistently emphasize detection performance, without advancing toward behavioral feedback or adaptive intervention frameworks.
Behavior Change Intervention
Previous studies have reviewed a variety of digital intervention strategies aimed at promoting healthy eating behaviors. Zhang et al. (2020) proposed a generative design framework for technology-mediated eating interventions, while Pan et al. (2021) surveyed digital tools, including mobile apps and wearable devices, designed to support healthy eating in occupational settings. However, these reviews primarily focus on organizing design considerations and intervention strategies, often relying on traditional education or standalone gamification without incorporating sensing technologies or perceptual computing. As a result, they typically lack integration with behavioral sensing or interactive feedback mechanisms.
In summary, while many existing reviews emphasize isolated stages such as intake detection, dietary assessment, or single-instance feedback, few focus on the complete behavioral closed-loop process from sensing to intervention.
This fragmentation highlights the urgent need for a systematic review that frames ingestion health interventions within a behavioral closed-loop paradigm, grounded in perceptual computing and spanning the full pipeline from sensing to feedback and intervention.
@@
To address these gaps, we present a structured review guided by the following research questions (RQs), aimed at evaluating how ingestion health interventions are conceptualized and implemented within behavioral closed-loop systems:
@@
  • We propose a behavioral closed-loop framework tailored to ingestive behavior interventions, and review the evaluation of effectiveness and user acceptance in current literature.
@@ 2.1. Paper Retrieval
Given that most relevant publications place greater emphasis on the feedback or intervention aspect, we also include studies that do not cover the full pipeline from sensing to intervention. After careful discussion, we decided to include all publications in which sensor-based interactive devices or perceptual computing methods are involved at any stage within the behavioral closed-loop process, ranging from sensing to feedback or intervention.
@@
  • Health Intervention: A system-generated feedback mechanism, whether active or passive, that aims to alter the user’s behavior with the goal of improving health from a medical perspective;
@@
Specifically, Benoit and Tracy (2008) defined ingestive behavior as ”a complex process controlled by multiple physiological and psychological mechanisms”, and divided it to two general categories, including ”behaviors involved in seeking, locating, and acquiring foods” and ”behaviors associated with the act of ingestion itself, including chewing and swallowing”. We refer to these two categories as ”appetitive behaviors” and ”consummatory behaviors”, and additionally added ”post-ingestive behaviors” to cover physiological feedback mechanisms and related behaviors after ingestive behavior. Moreover, Mendel and Wu (2010) described perceptual computing as a sensor-based method “associated with machinery for computing and reasoning with perceptions and data,” while Poole (2013) noted that “health interventions leverage technology to deliver timely support, monitoring, and guidance to influence health behaviors and outcomes.” These definitions align closely with the conceptual scope of our study. Accordingly, we adopted and refined the above terms to form the foundation of our inclusion criteria and analytical framework.
@@
{mdframed}
[backgroundcolor=gray!15] TF-IDF: eating, behavior, feedback, food, intake
Bi-gram: behavior(al) change, eating rate, eating habits, eating behavior, to improve
We selected three primary databases for our literature retrieval: the ACM Digital Library, IEEE Xplore, and Google Scholar. These sources are considered to comprehensively cover publications closely related to sensing technologies, interactive devices, and the Human–Computer Interaction (HCI) domain. Given the variation in keyword syntax and thesaurus constraints across platforms, we tailored the search strings and scoping targets for each database accordingly. The final keyword formulations and corresponding search results for each source are summarized below. All searches were conducted on April 4, 2025, without restricting the publication time interval.
{mdframed}
[backgroundcolor=gray!15] Keywords (combined with AND):
  • eating-related: eat*, ”food”, diet*, ”intake”
  • intervention-related: interven*, ”feedback”, ”behavior(al) change”, promot*, adjust*, design*, improv*, alter*
  • technology-related: ”wearable”, ”sensor(s)”, ”device(s)”, ”interactive”
ACM Digital Library
  • query target: Title, Keywords, Abstract
  • filters: No keynote, tutorial, invited talk, panel
  • inquiry result: 679
IEEE Explore
  • query target: Document Title (eating), Abstract (eating, intervention), All Metadata (technology)
  • filters: Conferences, Journals
  • inquiry result: 911
Google Scholar
  • filters: No books
  • inquiry result: 100
@@ 2.2. Screening
To further narrow down the set of retrieved articles to those aligned with the defined scope of our review, two authors conducted a structured two-stage screening process in accordance with PRISMA guidelines. During the initial screening, which involved title and abstract review followed by full-text assessment, each article was evaluated against a predefined set of inclusion and exclusion criteria. Articles were labeled as “include”, “exclude”, or “uncertain” and any disagreements were resolved through discussion until consensus was reached. The selection criteria used for this process are outlined below and have been adapted to follow PRISMA recommendations.
@@
Our overall multi-staged screening and selection process is depicted in Figure 2. The final set of review paper set includes 136 papers overall. Annual publication count is shown in Figure 3, illustrating a growing research interest in ingestion health in recent years. We identified 93 distinct publication venues across all included studies, with full details provided in LABEL:tab:venues_list. Among them, the following venues featured two or more included papers: ACM Conference on Human Factors in Computing Systems (CHI, n=16), ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT)/UbiComp (N=16), ACM Interaction Design and Children Conference (IDC, n=2), ACM Designing Interactive Systems Conference (DIS, n=4), IEEE Sensors (n=2), IEEE International Conference on Robot and Human Interactive Communication (RO-MAN, n=2), ACM/IEEE International Conference on Human-Robot Interaction (HRI, n=2), PervasiveHealth (n=2), SIGGRAPH Asia (n=2), International Conference on Tangible, Embedded, and Embodied Interaction (TEI, n=2), Innovations in Power and Advanced Computing Technologies (i-PACT, n=2), International Conference on Computing Communication and Networking Technologies (ICCCNT, n=2), World Congress on Medical Physics and Biomedical Engineering (n=2).
@@
To facilitate systematic analysis, we created a Google Sheets table with 57 columns. The column design and the taxonomy of core features (corresponding to Chapter 4 and Chapter 5) were iteratively developed. Two authors first defined preliminary categories during the filtering stage, followed by collaborative refinement of the taxonomy through a full review of the papers. Afterward, papers were split among authors for detailed annotation, with further refinements made throughout the reading process.
@@ 3. Result Overview
Following the review methodology and the research questions outlined above, this section presents an overview derived from the selected studies, including a classification based on Just-in-Time (JIT), and the primary behavioral closed-loop framework that synthesizes the key stages from sensing to intervention in ingestion health.
@@ 3.1. Primary Taxonomy based on JIT
Given the critical role of timing and context-awareness in optimizing the effectiveness of behavioral interventions, as highlighted in previous reviews (Hsu et al., 2025), we classified the included studies based on whether their intervention strategies exhibit JIT characteristics. With a primary overview based on JIT, this classification offers a high-level understanding of the integration mechanisms across existing ingestive behavior interventions, encompassing both sensing and feedback/intervention components.
JIT interventions are defined as approaches that deliver support dynamically at behaviorally or contextually opportune moments, relying on real-time assessments of users’ behavioral, physiological, or environmental states (Nahum-Shani et al., 2018; Sarker et al., 2014). Determining the appropriate timing for intervention typically involves identifying periods when users are vulnerable to adverse health outcomes and are physically, cognitively, and socially available to receive support (Choi et al., 2019; Sarker et al., 2014).
As shown in Figure 4(a), early researches in this domain primarily focused on non-JIT interventions, exhibiting a steady but slow growth over the years. Beginning from 2007, studies incorporating JIT intervention strategies began to emerge, and by 2011, the cumulative number of JIT publications surpassed that of non-JIT publications. A marked acceleration in the growth of JIT-related work can be observed around 2015–2016, which can be attributed to the rapid advancement of wearable sensing technologies and the increasing availability of real-time behavioral and contextual data. The distribution of JIT versus non-JIT studies is further illustrated in Figure 4(b), where JIT interventions account for a substantial majority (73.7%) of the included literature. Together, these findings provide a high-level understanding of the temporal trends and overall characteristics of research in the field of ingestion health interventions.
@@
3.2. Behavioral Closed-loop Framework
After providing an overview of the presence of JIT characteristics in the reviewed literature, we shift our focus to systematically examining behavior sensing and intervention paradigms, structured through a behavioral closed-loop framework. This framework serves as the analytical foundation for addressing the review questions outlined earlier, providing a coherent structure for understanding the current landscape of ingestive behavior interventions throughout the remainder of the paper.
In the broader field of human-computer interaction (HCI), ”closed-loop” architectures are defined as systems in which signals from the user’s brain and body are continuously collected, analyzed, and categorized as inputs to an adaptive controller, which then triggers appropriate responses at the interface (Fairclough, 2015). Such systems facilitate a continuous, bidirectional dialogue between the user and the technology. Furthermore, interaction loops in HCI can be conceptualized as exchanges between mutually interacting agents, where both the user and the system perceive and act upon each other’s states over time (Murray-Smith et al., 2024). In the context of ingestive behavior interventions discussed in this paper, we define ”behavioral closed-loop” as dynamic systems or approaches modeled as exchanges between interacting agents. In these systems, user ingestive behavior is monitored, interpreted, and utilized to deliver adaptive feedback or interventions, with subsequent behavioral changes tracked to iteratively refine future measures. This framework emphasizes the dynamic, responsive, and iterative nature of behavioral closed-loop interventions, capturing the essential stages involved in intervening from perception to ingestive behavior.
Specifically, the stages in behavioral closed-loop system include identifying target behaviors, sensing behaviorally relevant data, interpreting the observed phenomena through reasoning processes, developing feedback or intervention strategies, and monitoring subsequent behavioral adjustments to close the loop. This structure guides the subsequent analysis and provides a detailed understanding of how perceptual computing techniques are applied across different stages of intervention design. An overview of the primary behavioral closed-loop paradigm and its staged evolution is illustrated in Figure 5. While existing studies often focus on specific stages rather than achieving a fully closed-loop design, we structure the subsequent analysis and synthesis based on the ideal of a complete behavioral closed-loop framework.

### v4

# A Review of Behavioral Closed-Loop Paradigm from Sensing to Intervention for Ingestion Health

@@ Abstract.
Figure 1. Paradigm of ingestion health from sensing to intervention
Ingestive behavior plays a critical role in health, yet many existing interventions remain limited to static guidance or manual self-tracking. With the increasing integration of sensors, context-aware computing, and perceptual computing, recent systems have begun to support closed-loop interventions that dynamically sense user behavior and provide feedback during or around ingestion episodes. In this survey, we review 136 studies that leverage sensor-enabled or interaction-mediated approaches to influence ingestive behavior. We propose a behavioral closed-loop paradigm rooted in context-aware computing and inspired by HCI behavior change frameworks, comprising four components: target behaviors, sensing modalities, reasoning and intervention strategies. A taxonomy of sensing and intervention modalities is presented, organized along human- and environment-based dimensions. Our analysis also examines evaluation methods and design trends across different modality-behavior pairings. This review reveals prevailing patterns and critical gaps, offering design insights for future adaptive and context-aware ingestion health interventions.
@@ 1. Introduction
Ingestion health plays a critical role in shaping health outcomes, influencing quality of life, longevity, and overall well-being (Jovanov et al., 2014; Mesas et al., 2012). Despite its importance, current ingestion health interventions often rely on static tools such as food diaries (Thompson and Subar, 2017), self-report questionnaires (Shim et al., 2014), and periodic counseling sessions (Greaves et al., 2011), which place a heavy burden on users and provide limited behavioral feedback (Raber et al., 2021; Takemoto et al., 2018). As a result, they often fail to dynamically support individuals in developing sustainable and healthy ingestive behaviors over time.
The emergence of ubiquitous computing has long envisioned systems that seamlessly integrate into everyday life, sensing and responding to context to enhance human activity (Weiser, 1991). Advances in sensing technologies have further enabled low-burden, continuous monitoring of ingestive behaviors. In particular, multimodal sensor fusion has made unobtrusive ingestion monitoring feasible (Bell et al., 2020), allowing for the collection of rich behavioral data with minimal user effort.
Context-aware computing is a core capability of ubiquitous computing (Abowd and Mynatt, 2000), and it advocates modular architectures in which continuous sensing informs higher-level reasoning to enable adaptive interventions. Dey (2001) defined context-aware systems as those that ”use context to provide relevant services to the user, where relevancy depends on the user’s task”. Accordingly, sensing is not merely about acquiring raw data, but about capturing context to enable situation inference. Intervention, then, corresponds to the application component that, as Abowd and Mynatt (2000) emphasized, makes decisions and takes actions based on inferred context. These principles support personalized, adaptive interventions grounded in real-time behavioral context (Bell et al., 2020), marking a shift from static, user-driven methods to context-aware, behavior-informed systems.
Studies of ingestion health related to the aforementioned context have been systematically reviewed, with prior reviews primarily focusing on three key aspects: dietary assessment, wearable ingestion detection, and behavior change intervention.
i. Dietary Assessment: Skinner et al. (2020) and Wang et al. (2022) reviewed sensor-based frameworks for improving dietary assessment accuracy. However, their scope remains limited to intake monitoring, without addressing how such data can inform adaptive interventions.
ii. Wearable Ingestion Detection: Reviews by Bell et al. (2020) and He et al. (2020) systematically analyzed sensor types and detection pipelines, while others focused on chewing-specific sensing (Wei et al., 2022; Mortazavi and Gutierrez-Osuna, 2023). These works emphasize detection performance but rarely consider downstream behavior change frameworks.
iii. Behavior Change Intervention: Zhang et al. (2020) and Pan et al. (2021) reviewed digital tools for eating-related interventions but primarily centered on static strategies such as education or gamification, with limited integration of sensing or context-aware intervention.
In summary, prior reviews tend to isolate sensing, reasoning, or intervention, without linking them into an integrated behavior change pipeline. This highlights the need for a systematic review that explicitly integrates ingestion health interventions and perceptual computing into a cohesive system. To address this, we propose a behavioral closed-loop paradigm for ingestion health that encompasses sensing, reasoning, and intervention. This paradigm builds on Salber’s Context Toolkit (Salber et al., 1999), inspired by Personal Informatics Framework (Li et al., 2010) and Digital Behavior Change Intervention (Hekler et al., 2016), aiming to bridge the scattered components identified in previous reviews. Moreover, it advances from passive context display toward more active and adaptive forms of behavioral scaffolding.
@@
Building on this conceptual foundation, we conduct a structured review guided by the following research questions (RQs), aimed at evaluating how ingestion health interventions are conceptualized and implemented within behavioral closed-loop systems:
@@
  • We propose a behavioral closed-loop paradigm, grounded in context-aware computing and informed by HCI behavior change frameworks, to guide ingestive behavior interventions, and review the evaluation of effectiveness and user acceptance in current literature.
@@ 2.1. Paper Retrieval
Given that most relevant publications place greater emphasis on the intervention aspect, we also include studies that do not cover the full pipeline from sensing to intervention. After careful discussion, we decided to include all publications in which sensor-based interactive devices or perceptual computing methods are involved at any stage within the behavioral closed-loop process, ranging from sensing to intervention.
@@
  • Health Intervention: A system-generated mechanism aimed at altering user behavior for health improvement, categorized as passive interventions — providing feedback to foster self-awareness, and active interventions — delivering structured guidance or direct mechanisms to proactively guide behavior.
@@
While the paradigm is conceptually grounded in context-aware computing, we adopt ”perceptual computing” as a core operational term. This term provides a more implementation-oriented perspective on behavioral data processing in closed-loop systems building on context-aware principles. It focuses on gaps in prior reviews where perception and intervention were often treated as disconnected components.
Specifically, Benoit and Tracy (2008) defined ingestive behavior as ”a complex process controlled by multiple physiological and psychological mechanisms”, and divided it into two general categories, including ”behaviors involved in seeking, locating, and acquiring foods” and ”behaviors associated with the act of ingestion itself, including chewing and swallowing”. We refer to these two categories as ”appetitive behaviors” and ”consummatory behaviors”, and additionally added ”post-ingestive behaviors” to cover physiological feedback mechanisms and related behaviors after ingestive behavior. Moreover, Mendel and Wu (2010) described perceptual computing as a sensor-based method ”associated with machinery for computing and reasoning with perceptions and data”, while Poole (2013) noted that ”health interventions leverage technology to deliver timely support, monitoring, and guidance to influence health behaviors and outcomes”. These definitions align closely with the conceptual scope of our study. Accordingly, we adopted and refined the above terms to form the foundation of our inclusion criteria and analytical framework.
@@
TF-IDF: eating, behavior, feedback, food, intake Bi-gram: behavior(al) change, eating rate, eating habits, eating behavior, to improve
We selected three primary databases for our literature retrieval: the ACM Digital Library, IEEE Xplore, and Google Scholar. These sources are considered to comprehensively cover publications closely related to sensing technologies, interactive devices, and the Human–Computer Interaction (HCI) domain. Given the variation in keyword syntax and thesaurus constraints across platforms, we tailored the search strings and scoping targets for each database accordingly. The final keyword formulations are summarized below, with corresponding search details for each source shown in Appendix A. All searches were conducted on April 4, 2025, without restricting the publication time interval.
Keywords (combined with AND): eating-related: eat*, ”food”, diet*, ”intake” intervention-related: interven*, ”feedback”, ”behavior(al) change”, promot*, adjust*, design*, improv*, alter* technology-related: ”wearable”, ”sensor(s)”, ”device(s)”, ”interactive”
@@ 2.2. Screening
To further narrow down the set of retrieved articles to those aligned with the defined scope of our review, two authors conducted a structured two-stage screening process with PRISMA guidelines. During the initial screening, which involved title and abstract review followed by full-text assessment, each article was evaluated against a predefined set of inclusion and exclusion criteria. Articles were labeled as ”include”, ”exclude”, or ”uncertain” and any disagreements were resolved through discussion until consensus was reached. The selection criteria used for this process are outlined below and have been adapted to follow PRISMA recommendations.
@@
Our overall multi-staged screening and selection process is depicted in Figure 2. The final set of review paper set includes 136 papers overall. Annual publication count is shown in Figure 3, illustrating a growing research interest in ingestion health in recent years. We identified 93 distinct publication venues across all included studies, with full details provided in Table LABEL:tab:venues_list. Among them, the following venues featured two or more included papers: ACM Conference on Human Factors in Computing Systems (CHI, n=16), ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT)/UbiComp (N=16), ACM Interaction Design and Children Conference (IDC, n=2), ACM Designing Interactive Systems Conference (DIS, n=4), IEEE Sensors (n=2), IEEE International Conference on Robot and Human Interactive Communication (RO-MAN, n=2), ACM/IEEE International Conference on Human-Robot Interaction (HRI, n=2), PervasiveHealth (n=2), SIGGRAPH Asia (n=2), International Conference on Tangible, Embedded, and Embodied Interaction (TEI, n=2), Innovations in Power and Advanced Computing Technologies (i-PACT, n=2), International Conference on Computing Communication and Networking Technologies (ICCCNT, n=2), World Congress on Medical Physics and Biomedical Engineering (n=2).
@@
To facilitate systematic analysis, we created a Google Sheets table with 57 columns. The column structure was iteratively developed to support the discovery and refinement of the taxonomy of core features (corresponding to Chapter 4 and Chapter 5). Two authors first defined preliminary categories during the filtering stage, followed by collaborative refinement of the taxonomy through a full review of the papers. Afterward, papers were split among authors for detailed annotation, with further refinements made throughout the reading process.
@@ 3. Result Overview
Following the review methodology and research questions outlined above, this section presents a two-part overview of papers selected in Section 2.2. We first classify the publications based on Just-in-Time (JIT) characteristics, highlighting their temporal and contextual responsiveness. Then, we introduce our behavioral closed-loop paradigm grounded in context-aware computing and inspired by HCI behavior change frameworks. This paradigm outlines key stages from sensing to intervention and serves as the analytical foundation for the subsequent taxonomy and discussion.
@@ 3.1. Primary Taxonomy based on JIT
Building on prior work emphasizing the importance of timing and context-awareness in behavioral interventions (Hsu et al., 2025), we classified the included studies based on whether their intervention strategies exhibit JIT characteristics. This classification offers a high-level understanding of the integration mechanisms across existing ingestion health systems, encompassing both sensing and intervention components.
JIT interventions are defined as approaches that deliver support dynamically at behaviorally or contextually opportune moments, determined by real-time assessments of users’ behavioral, physiological, or environmental states—particularly when individuals are vulnerable and receptive to intervention (Nahum-Shani et al., 2018; Sarker et al., 2014; Choi et al., 2019). As illustrated in Figure 4(a), JIT interventions have gained increasing attention since their emergence around 2007, surpassing non-JIT studies by 2011. A notable surge occurred after 2015, driven by advances in wearable sensing and the availability of real-time contextual data. Overall, JIT interventions account for 73.7% of the reviewed studies, as shown in Figure 4(b), highlighting their central role in contemporary ingestion health intervention design.
@@
3.2. Behavioral Closed-loop Paradigm
We propose a behavioral closed-loop paradigm that integrates various components — sensing, reasoning, and intervention — into a cohesive system. This paradigm is conceptually grounded in Human-Computer Interaction (HCI), drawing inspiration from context-aware computing and behavior change fields. Specifically, context-aware theories provide the theoretical foundation for system components, while HCI behavior change frameworks inform the behavioral logic of the closed-loop paradigm, resulting in an integrated closed-loop system.
3.2.1. Grounded in Context-Aware Computing
Context-aware computing emphasizes how interactive systems continuously sense, interpret, and respond to contextual information to enable adaptive, task-relevant interventions. Building on Dey’s foundational definition, context is defined as ”any information that can be used to characterize the situation of an entity”, and a context-aware system ”uses context to provide relevant information and/or services to the user, where relevancy depends on the user’s task(Dey, 2001). These two definitions align with our paradigm’s components in ingestion health systems: Sensing and Reasoning. Additionally, the Context Toolkit proposed by Salber et al. (1999) highlights a key principle: the separation of sensing, context interpretation, and application logic into modular components. Consequently, we have derived three main components in our paradigm:
  • Sensing: Raw sensor inputs are first collected as behavioral signals.
  • Reasoning: Contextual signals are combined and interpreted to infer the user’s situation.
  • Intervention: Application components make task-relevant decisions and deliver adaptive actions based on interpreted context.
3.2.2. Inspired from Behavior Change Frameworks
We then turn to behavioral change frameworks for a process-oriented perspective that clarifies how components can be organized temporally and functionally. The Personal Informatics Framework proposed by Li et al. (2010) describes a five-stage cyclical process: preparation, collection, integration, reflection, and action. This framework inspired us to conceptualize our paradigm as a behavioral loop, mapping collection to Sensing, integration to Reasoning, and the reflection–action sequence to Intervention, framing the paradigm as a loop that enables continuous adjustment. Building on this, the Digital Behavior Change Interventions (DBCI) framework (Hekler et al., 2016) provides system-level insights by integrating sensing, tailoring, action, and interaction into a unified architecture for adaptive interventions (Yardley et al., 2016). This approach led us to design the components to function not as isolated modules but as an integrated system.
While our paradigm focuses on three system-level components, it draws inspiration from the five-stage PIF process. Some stages, such as preparation, are embedded implicitly within broader sensing or reasoning phases.
3.2.3. Strengthened with Closed-Loop Concept
To reinforce the ”loop” structure, we draw from broader ”closed-loop” architectures in HCI. According to Fairclough (2015), closed-loop systems continuously monitor user states and feed those signals into adaptive controllers to trigger context-appropriate responses. Murray-Smith et al. (2024) further conceptualize such systems as mutually interactive agents, where both the user and the system perceive and act on each other’s changing states over time.
Drawing from these perspectives, we develop our behavioral closed-loop paradigm as a dynamic, sensing-driven system that continuously interprets user behavior and delivers adaptive interventions, enabling iterative feedback over time and context. This feedback can arise from real-time behavioral data, user-state inference, or performance evaluation, enabling adaptive recalibration across the loop. Instead of a monolithic process, this paradigm comprises recurring stages: identifying target behaviors, sensing relevant data, delivering interventions, and evaluating outcomes to guide future actions. When applied across everyday environments, these loops form dense, time- and context-sensitive patterns, realizing the vision of ubiquitous computing through distributed, context-aware interventions. Figure 5 illustrates the core structure and evolution of the behavioral closed-loop paradigm, informing our subsequent taxonomy and system-level analysis.